Peter leaves at a speed of 16km/h and 1 1/2 hours later Bob starts out at 20 km/h. How long does it take until Bob has covered 1/2 Peter's distance? What distance will each have travelled at this point?
Answer by ankor@dixie-net.com(22740)

Peter leaves at a speed of 16km/h and 1 1/2 hours later Bob starts out at 20 km/h.
How long does it take until Bob has covered 1/2 Peter's distance?
:
Let t = Bob's travel time (in hrs)
then
(t+1.5) = Peter's travel time
:
Write a distance equation dist = speed * time
:
P's dist = twice B's dist
16(t+1.5) = 2(20t)
16t + 24 = 40t
24 = 40t - 16t
24t = 24
t = 1 hr for Bob to cover half Peter's distance
:
What distance will each have traveled at this point?
P's dist: 16(1+1.5) = 40 km
B's dist: 20(1) = 20 km